OpenStudy (anonymous):

find the number c that satisfies the conclusion of the Mean Value Theorem for f on the closed interval [1,3]

OpenStudy (anonymous):

for what equation ?

OpenStudy (anonymous):

depends largely on \(f\)

OpenStudy (anonymous):

F(x)= x^3-7x+6

OpenStudy (anonymous):

srios can you help me?

OpenStudy (anonymous):

\[f(3)=3^3-7\times 3+6=27-21+6=12\] \[f(1)=1-7+6=0\] \[f(3)-f(1)=12\] \[\frac{f(3)-f(1)}{3-1}=\frac{12}{2}=6\]

OpenStudy (anonymous):

\(f'(x)=3x^2-7\) set \[3x^2-7=6\] and solve for \(x\)

OpenStudy (anonymous):

Thank you!